max is building a rectangular planter without a top. the planter will be 5 inches wide, 15 inches long, and 8 inches tall. how much wood is needed to make the planter?
Found 2 solutions by ikleyn, linhamnor:
Answer by ikleyn(52781)

.

Calculate the area of each of four vertical faces.

Calculate the area of the bottom.

Keep in mind that the area of a rectangle is the product of its dimensions.

Having these instructions, boldly go forward.


I hope you are competent in making arithmetic operations, such as multiplying and adding integer numbers,
because these skills are prerequisites to this problem. So, I hope that there is no need to teach you
to add and to multiply moderate integer numbers.


Making these simple calculations, you have a great chance and wonderful opportunity
to develop your skills further - up to heavenly heights.


Happy calculations  ( ! )



Answer by linhamnor(1)

5 vertices
base = L x W = 15x5 = 75
long sides = 2(LxW) = 2(15x8) = 240
short sides = 2(WxH) = 2(5x8) = 40
75 + 240 + 40 = 395 inches squared